OpenStudy (anonymous):

Find functions f and g so that h(x) = (f ∘ g)(x). h(x) = (6x - 14)^8

OpenStudy (anonymous):

do u know this question has many answers

OpenStudy (anonymous):

Oh, really? Here's the options I was given- A. f(x) = 6x - 14, g(x) = x8 B. f(x) = 6x8 - 14, g(x) = -14 C. f(x) = x8, g(x) = 6x - 14 D. f(x) = (6x)8, g(x) = -14

OpenStudy (anonymous):

f(x) could be f(x) =x^8, and g(x) =6x-14

OpenStudy (anonymous):

Oh, wow, thank you so much! I appreciate it!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

for example

OpenStudy (anonymous):

f(x) =x^4 and g(x) (6x-14)^2

OpenStudy (anonymous):

Would that also be correct?

OpenStudy (anonymous):

oh u have multiple choice

OpenStudy (anonymous):

i just saw it now

OpenStudy (anonymous):

stick to the choices u have

OpenStudy (anonymous):

but yes if there was no multiple choice u would have a lot of answer to that question

OpenStudy (anonymous):

for example

OpenStudy (anonymous):

f(x) = x, and g(x)=(6x-14)^8

OpenStudy (anonymous):

Egad, thank goodness for multiple choice questions. Thank you so much.

OpenStudy (anonymous):

ur more than welcome, but did u know how this worked?

OpenStudy (anonymous):

I sort of think so? If you have the time to expand it, I'd appreciate it.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

composite functions are very important in calculus u need really be able to visualize the inside out, especially for the sake of the chain rule

OpenStudy (anonymous):

or another term it may be called function of a function

OpenStudy (anonymous):

so what does f o g mean?

OpenStudy (anonymous):

it means everything in g throw it in the bin of f

OpenStudy (anonymous):

so let f(x) be ur container

OpenStudy (anonymous):

Ok, so far so good...

OpenStudy (anonymous):

so try this one f(x) = x^x^x, and g(x) = e^x^2 find g o f, and f o g

OpenStudy (anonymous):

if you solve this problem then u can solve any composite function

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

Sorry, trying to work it out.

OpenStudy (anonymous):

ok take ur time give me a shout when u r done

OpenStudy (anonymous):

So f o g would be... g(x) = e^x^2 = X^x[(e^x^2)]^x Oh, wait, I think I'm lost.

OpenStudy (anonymous):

u need to throw e^x^2 in every x

OpenStudy (anonymous):

Oooooh.

OpenStudy (anonymous):

squeeze ur mind and focus this will give u a deep understanding, even if u don't get the answer

OpenStudy (anonymous):

this is my point u need to think in depth

OpenStudy (anonymous):

ok i will leave now u take ur time am sure u can solve, then give me the answer anytime u want ok